OpenStudy (anonymous):

solve cosx=sin2x on [0,2pi]

jimthompson5910 (jim_thompson5910):

cos(x)=sin(2x) cos(x)=2*sin(x)*cos(x) cos(x)-2*sin(x)*cos(x) = 0 cos(x)( 1-2*sin(x) ) = 0 I'll let you finish

OpenStudy (anonymous):

I got 1/2=sin(x) so x=\[\pi/6, 5\pi/6\]

jimthompson5910 (jim_thompson5910):

that's part of the solution set

OpenStudy (anonymous):

Oh! Okay, thanks a bunch!

jimthompson5910 (jim_thompson5910):

what's the rest of the solution set

OpenStudy (anonymous):

\[11\pi/6, 7\pi/6\] ? is that right or am I not getting this at all?

jimthompson5910 (jim_thompson5910):

no idk how you got that

jimthompson5910 (jim_thompson5910):

if cos(x)( 1-2*sin(x) ) = 0 then cos(x) = 0 or 1 - 2sin(x) = 0

jimthompson5910 (jim_thompson5910):

you already solved 1 - 2sin(x) = 0, so just solve the first equation to get your two other solutions

OpenStudy (anonymous):

alright, so it would be \[\pi/2, 3\pi/2\]

jimthompson5910 (jim_thompson5910):

yep, those are added to the other two solutions you got

jimthompson5910 (jim_thompson5910):

so your four solutions are pi/6, 5pi/6, pi/2, 3pi/2

OpenStudy (anonymous):

Thanks a bunch for the help.

jimthompson5910 (jim_thompson5910):

you stop at these 4 because you're restricted to the interval [0, 2pi)

jimthompson5910 (jim_thompson5910):

yw